Solve the initial-value problem


1. y" - 3y' - 10y = 0, y(0) = 0, y'(0) = 7.


2. y"+14y' + 50y = 0,y(0) = 2, y'(0) = -17


3. 6y"-y'-y = 0, y(0) = 10, y'(0) = 0


4. 6y"+y'-y = 0, y(0) = -1, y'(0) = 3


1
Expert's answer
2021-05-24T12:32:50-0400

1.

"k^2-3k-10=0"

"k=\frac{3\pm \sqrt{9+40}}{2}"

"k_1=-2,k_2=5"

"y=c_1e^{-2x}+c_2e^{5x}"

"0=c_1+c_2"

"y'=-2c_1e^{-2x}+5c_2e^{5x}"

"7=-2c_1+5c_2"

"2c_2+5c_2=7"

"c_2=1,c_1=-1"

"y=-e^{-2x}+e^{5x}"


2.

"k^2+14k+50=0"

"k=\frac{-14\pm \sqrt{196-200}}{2}"

"k_{1,2}=\frac{-14\pm 2i}{2}=-7\pm i"

"y=e^{-7x}(c_1cosx+c_2sinx)"

"2=c_1"

"y'=-7e^{-7x}(c_1cosx+c_2sinx)+e^{-7x}(c_2cosx-c_1sinx)"

"-17=-7c_1+c_2"

"c_2=-3"

"y=e^{-7x}(2cosx-3sinx)"


3.

"6k^2-k-1=0"

"k=\frac{1\pm \sqrt{1+24}}{12}"

"k_1=-\frac{1}{3},k_2=\frac{1}{2}"

"y=c_1e^{-x\/3}+c_2e^{x\/2}"

"10=c_1+c_2"

"y'=-\frac{c_1e^{-x/3}}{3}+\frac{c_2e^{x/2}}{2}"

"0=-c_1\/3+c_2\/2"

"\frac{c_2-10}{3}+\frac{c_2}{2}=0"

"c_2=4,c_1=6"

"y=6e^{-x\/3}+4e^{x\/2}"


4.

"6k^2+k-1=0"

"k=\frac{-1\pm \sqrt{1+24}}{12}"

"k_1=-\frac{1}{2},k_2=\frac{1}{3}"

"y=c_1e^{-x\/2}+c_2e^{x\/3}"

"-1=c_1+c_2"

"y'=-\frac{c_1e^{-x/2}}{2}+\frac{c_2e^{x/3}}{3}"

"3=-\frac{c_1}{2}+\frac{c_2}{3}"

"3(c_2+1)+2c_2=18"

"c_2=3,c_1=-4"

"y=-4e^{-x\/2}+3e^{x\/3}"